OpenStudy (anonymous):

Question 1.1. Which property is demonstrated? 5 + 3w + 2 = 5 + 2 + 3w (Points : 1) commutative property of addition commutative property of multiplication associative property of addition associative property of multiplication

OpenStudy (anonymous):

commutative of addition. a+b=b+a

OpenStudy (anonymous):

it is associative property of addition

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

associative property of addition (a+b)+c=a+(b+c)